A man is 24 years older than his son. Three years ago he was eight years less than twice his son's age three years from now. How old is the man?
Answer by stanbon(75887)

A man is 24 years older than his son. Three years ago he was eight years less than twice his son's age three years from now. How old is the man?
-------------
Equations:
f = s + 24
f-3 = 2(s+3)-8
-----
Substitute for "f" and solve for "s"::
s+24-3 = 2s - 2
s+21 = 2s-2
s = 23 yrs (son's age now)
---
f = s + 24 = 47 yrs (father's age now)
------------
Cheers,
Stan H.
---------------